OpenStudy (anonymous):

What is the arc length of the subtending arc for an angle of 72 degrees on a circle of radius 4?

OpenStudy (fortytherapper):

Arc Length: \[\frac{ \theta }{ 360 }(2\pi r)\] We were given: \[r = 4\] \[\theta = 72\] Now we just plug and chug

likeabossssssss (likeabossssssss):

welcome to os